I have to find the LCM OF Y/10X^2 AND X/15Y^3. I am getting an answer of 30x^2y^3 and another answer to the same problem of 1/150xy^2. Can you show me the steps on how to do this problem correctly.This problem is not in the book. It is a handout. Thankyou!

Laurie
hubbsurg206@aol.com

Answer by edjones(8007)

The LCM of the the denominators is the same as the LCD.
10=2*5
15=3*5
the prime multiples in both are 2*3*5=30
The LCM= 30x^2y^3
.
Ed